# Superresolution for two incoherent optical sources with arbitrary intensities in two dimensions

### v1

@@ Abstract
The Rayleigh criterion has long served as a fundamental limit for the resolution of classical optical imaging. However, recent advances in quantum metrology have led to the quantum superresolution technique that can break Rayleigh’s curse and estimate the separation between a pair of incoherent point sources with nonvanishing precision. For two-dimensional optical systems, the quantum superresolution has been studied for the orthogonal components of separation between two incoherent equal-intensity point sources in a given coordinate system. However, the precision limit of estimating the whole separation, i.e., the distance, between two point sources remains unknown so far. In this paper, we investigate the estimation precision of the distance between two incoherent point sources with arbitrary intensities in a two-dimensional imaging system. Through the multiparameter quantum estimation theory, we obtain the ultimate estimation precision of the distance and show that it remains nonzero when the distance approaches zero, which surpasses the Rayleigh criterion. We further show that the precision can be enhanced by aligning the two point sources along specific directions if the point-spread functions of the two sources are not circularly symmetric, and find the optimial relative azimuth between the two point sources and the highest estimation precision of the distance. In addition to the distance estimation, we also consider the quantum estimation of the relative azimuth between two incoherent point sources. A surprising result is that the precision limit of the azimuth decays quadratically with the distance, which suggests that the azimuth cannot be resolved when the two point sources get sufficiently close to each other and is therefore inaccessible by the quantum superresolution technique in this case. This reveals a new and fundamental limitation on the resolvability of two incoherent point sources in multi-dimensional quantum imaging which cannot be addressed by optimizing the quantum measurements.
@@ I Introduction
A notable example is the longstanding challenge of resolving closely spaced optical sources, which is constrained by Rayleigh’s criterion in classical optics [11]. This criterion sets a diffraction-limited resolution threshold at $d=1.22\lambda/(2{\rm NA})$, with $\lambda$ the wavelength of the light and $\rm NA$} the numerical aperture of the imaging system. To overcome this limit, various approaches have been proposed. One straightforward method is to reduce the wavelength or increase the numerical aperture [12]. More fundamentally, techniques such as microscopy structured illumination [13, 14], photoactivatable localization [15] and stochastic optical reconstruction microscopy [16] have been proposed to surpass the Rayleigh limit by exploiting nonlinear optical responses, stochastic activation of fluorescent emitters, and computational reconstruction algorithms. While these techniques have achieved remarkable progress in imaging resolution, they typically require specialized fluorophores, sophisticated control of activation cycles, or complex image processing [17].
From a statistical perspective, when two incoherent optical sources have significant overlap on the image plane, the precision of estimating their separation is fundamentally limited by the Cramér-Rao lower bound [18, 19]. It follows that the estimation precision of the separation between two optical sources vanishes as the separation approaches zero by direct imaging. This statistical phenomenon is known as Rayleigh’s curse. However, Tsang et al. [20] attacked this problem from the perspective of quantum multiparameter estimation and surprisingly found that the quantum Cramér-Rao bound (QCRB) [21] for the separation of two incoherent point sources does not vanish given arbitrarily small separations if the quantum measurements on the point sources are optimized, which addresses Rayleigh’s curse.
@@
In the original superresolution protocol [20], the superresolution was proposed for two optical sources with one-dimensional point-spread functions . However, the point-spread functions of practical optical sources are essentially multidimensional, where the spatial structures of the point-spread functions become crucial to the imaging resolution. In this broader context, multiparameter quantum estimation theory provides a powerful tool to improve the imaging precision [58, 59, 60, 61]. Building upon this foundation, subsequent researches have extended the quantum superresolution to the estimation of components of the separation in orthogonal directions for two-dimsional point sources [62] and for three-dimsional point sources [63, 64, 65], and to more complex imaging tasks such as surface metrology [66], using balanced homodyne detection [67], and estimating the sizes of uniform line and disk sources [68], etc.
In this work, we investigate the quantum precision limits of two-dimensional superresolution for a pair of incoherent point sources with arbitrary relative intensities. Rather than estimating the Cartesian components of the separation individually, we focus on the quantum estimation of the whole separation, i.e., the distance, between two point sources. We show that, by optimizing the measurement basis, the quantum Fisher information for the distance remains finite when the two sources get close to each other, which breaks the Rayleigh limit. Furthermore, when the point-spread functions (PSF) of the two point sources are not circularly symmetric, the estimation precision exhibits a nontrivial dependence on the relative orientation between the two point sources. This orientation dependence enables further precision enhancement by aligning the two point sources along specific directions. We optimize the quantum Fisher information of the distance over the azimuth between the two point sources, and obtain the optimal azimuth and the maximum quantum Fisher inforation for the estimation of source distance. Taking Gaussian PSFs as an example, we find that the optimal orientation corresponds to the minor axis of the PSF. We further consider the quantum estimation for the azimuth itself, and surprisingly find that it decays quadratically with the distance between the two sources and vanishes in the limit of zero distance. This suggests that while quantum metrology can circumvent Rayleigh’s curse for the estimation of the distance between two point sources, it provides no information about the relative direction between the sources when they are sufficiently close to each other. This imposes a new limitation on the imaging resolution within the framework of quantum mechanics.
@@ II.1 Model of optical sources
This question has been addressed for one-dimensional imaging systems by the pioneering work of Tsang et al., and also been considered for the estimation of the orthogonal components of the separation between two point sources in two-dimensional imaging systems, as reviewed in the introduction section. In this work, we will focus on the estimation of the whole separation, i.e., the distance, as well as the relative orientation, between two point sources in two-dimensional imaging systems.
Figure 1: Two point optical sources located at $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$. The distance between the two sources is $r$, and the relative orientation between the two point sources is characterized by the azmuth $\alpha$, i.e., the angle between the displacement of the two sources and a given reference axis which is the $x$ axis in the figure.
@@ III.1 Quantum precision limit for separation of point sources
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{{\hbox{N_{{\rm tot}}^{}}}}\left(1-\epsilon^{2}\right)\left(\kappa_{x}\kappa_{y}-\eta^{2}\right)}{\kappa_{r^{\perp}}},$ (20)
@@
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{{\hbox{N_{{\rm tot}}^{}}}}\left(\kappa_{x}\kappa_{y}-\eta^{2}\right)}{\kappa_{r^{\perp}}}.$ (22)
Any deviation from this balance leads to a degradation in precision. Counterintuitively, this suggests that introducing asymmetry to the intensities of two sources, while increasing the visual contrast of the intensity distribution on the image plane, actually reduces the resolvability of the two sources. Physically, this is because the resolution of two optical sources relies on the distingsuishability between them and lowering the intensity of either source will decrease its distinguishability from the other. For example, if the intensity of one source is tuned to zero, the visual contrast on the image plane reaches the maximum, but that source can never be distinguished from the other as it is impossible to determine its location when its intensity is zero.
@@ III.2 Optimization for precision limit of distance
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{{\hbox{N_{{\rm tot}}^{}}}}\kappa_{x}\kappa_{y}}{\sin^{2}\alpha\kappa_{x}+\cos^{2}\alpha\kappa_{y}},$ (27)
@@ III.3 Example: two-dimensional Gaussian point-spread functions
Figure 3: Density plot of the distance precision $\mathcal{H}_{r}$ for two Gaussian point-spread functions with respect to the azimuth $\alpha$ and and the raito $\epsilon$ of the photon number difference to the total photon number as $r\rightarrow 0$. $\sigma_{1}=1$, $\sigma_{2}=1.4$ and $\beta=0.4$. It can be observed that when the azimuth $\alpha$ is fixed, the estimation precision of the distance $r$ will decrease as the ratio $\epsilon$ increases and the optimal estimation precision can be always reached when the photon numbers of the two sources are the same, i.e., $\epsilon=0$, showing the negative impact of photon number difference on the superresolution precision. On the other hand, when $\epsilon$ is fixed, the estimation precision of the distance varies with the azimuth $\alpha$, and the maximum precision of the distance $r$ can be obtained when $\alpha$ is approximately $\pi/7$, which corresponds to the direction along the minor axis of the Gaussian point-spread function, implying the precision of estimating the distance can be enhanced by optimizing the relative orientation between the two point sources.
@@
which implies that it is always possible to increase the superresolution precision by aligning the two point sources whenever the Gaussian point-spread function is not circularly symmetric. On the other hand, when the Gaussian point-spread function has circular symmetry, the overlap betweent the two point-spread functions always remains the same regardless of the aligning of the two sources, so in this case, the precision becomes orientation-independent and there is no advantage in the precision by aligning the two sources.
@@
To illustrate the above results numerically, Fig. 3 plots the estimation precision $\mathcal{H}_{r}$ with respect to $\alpha$ and $\epsilon$ in the limit $r\rightarrow 0$. It can be seen that for a fixed azimuth $\alpha$, the superresolution precision $\mathcal{H}_{r}$ is maximized at $\epsilon=0$, corresponding to the case of equal photon intensities. As $\epsilon$ increases, $\mathcal{H}_{r}$ decreases and eventually vanishes with strong photon intensity imbalance. Moreover, for a given $\epsilon$, $\mathcal{H}_{r}$ shows clear dependence on the azimuth $\alpha$ between the two point sources, reaching its maximum at $\alpha_{{\rm opt}}=\tan^{-1}\frac{\sigma_{2}^{2}-\sigma_{1}^{2}-\Re}{2\beta\sigma_{1}\sigma_{2}}$, which coincides with the minor axis of the Gaussian point-spread function.
IV Estimation precision of azimuth
For two-dimensional imaging systems, not only the distance between two point sources is unknown, the relative orientation between the two sources is not known either. For the task of imaging, determining the orientation between two point sources is also critical, so in the following, we will study the estimation precision for the azimuth between the two sources, i.e., the angle between the displacement of the two sources and a reference axis.
The precision limit of estimating the azimuth $\alpha$ can be derived by the quantum Cramér-Rao bound, similar as the estimation of the distance $r$. As the derivation is lengthy, it is presented in Appendix C, which considers arbitrary photon intensities for the two sources. When the distance $r$ between the two sources is small, i.e., $r\ll 1$, the precision limit of estimating $\alpha$ can be further simplified to
$\mathcal{H}_{\alpha}=\frac{r^{2}N_{{\rm tot}}\left(1-\epsilon^{2}\right)\left(\kappa_{x}\kappa_{y}-\eta^{2}\right)}{\kappa_{r}}+O(r^{4}).$ (47)
While this result looks simple at the first glance, it shows a surprising quadratic scaling in $r$, implying that in the limit $r\rightarrow 0$, the quantum estimation precision $\mathcal{H}_{\alpha}$ of the azimuth $\alpha$ vanishes, so no information about $\alpha$ can be extracted in this case. This shows a new and fundamental limitation on quantum imaging, that is the inaccessibility of the azimuth $\alpha$ in the limit $r\rightarrow 0$ with any finite photon numbers, even if the quantum measurement is optimized!
Fortunately, when $r$ is exactly equal to zero, the azimuth $\alpha$ doest not make any sense as the two point sources coincide with each other, so it is actually not necessary to estimate $\alpha$ in this case. When the distance $r$ is small but not zero, it is still possible to circumvent the above issue by increasing $N$ to the order of $O(r^{-2})$ so that the precision $\mathcal{H}_{\alpha}$ is of order $O(1)$. This indicates that a quadratically increasing resource of photons is necessary to keep a finite precision of estimating the azimuth $\alpha$, which imposes a condition on the number of photons that one needs to allocate in order to determine the azimuth $\alpha$ by quantum measurement.
When the two point-spread functions are Gaussian, the estimation precision of the azimuth $\alpha$ becomes
$\mathcal{H}_{\alpha}=\frac{r^{2}\left(1-\epsilon^{2}\right)N_{{\rm tot}}}{4\left[\sigma_{1}^{2}\sin^{2}\alpha+\sigma_{2}^{2}\cos^{2}\alpha-\beta\sigma_{2}\sigma_{1}\sin(2\alpha)\right]}+O(r^{4}).$ (48)
To visualize this dependence, Fig. 4 plots the precision $\mathcal{H}_{\alpha}$ with respect to the azimuth $\alpha$ and the distance $r$. For any fixed $\alpha$, the estimation precision $\mathcal{H}_{\alpha}$ vanishes at $r=0$, indicating that no information about $\alpha$ can be extracted in this case. As $r$ increases, $\mathcal{H}_{\alpha}$ grows. Moreover, for a given $r$, $\mathcal{H}_{\alpha}$ exhibits a clear periodic dependence on the azimuth $\alpha$ between the two point sources.
Figure 4: Contour plot of the precision $\mathcal{H}_{\alpha}$ for two Gaussian point-spread functions as a function of the distance $r$ and the azimuth $\alpha$, where $\epsilon=0.4$ $\sigma_{1}=1$, $\sigma_{2}=1.4$ and $\beta=0.4$. The figure shows that for any fixed $\alpha$, the estimation precision $\mathcal{H}_{\alpha}$ vanishes as $r\rightarrow 0$.
V Conclusion
@@
Moreover, we found that the superresolution precision for distance estimation can be significantly enhanced by aligning the sources along specific directions, and derived the optimal relative azimuth and the highest estimation precision of the distance between the two sources. This orientation dependence opens a new avenue for optimizing imaging strategies in practical systems with non-circular symmetry.
However, our analysis also reveals a new fundamental limitation on quantum imaging: while the information of distance remains accessible, the estimation precision for the azimuth between two point sources drops quadratically as the distance decreases. This indicates that the relative direction between two spaced sources becomes inherently unresolvable in the quantum regime when they are sufficiently close to each other, revealing an intrinsic limit of quantum imaging that cannot be mitigated through measurement optimization.
In summary, our findings refine the understanding of quantum superresolution and its ultimate capabilities and limitations, offering both theoretical insight and practical guidance for the application of quantum superresolution in realistic scenarios.
@@ Appendix A Resolution limit of direct imaging
$\mathcal{J}_{ij}=\int_{-\infty}^{\infty}dy\int_{-\infty}^{\infty}dx\frac{1}{\Lambda(x,y)}\frac{\partial\Lambda(x,y)}{\partial g_{i}}\frac{\partial\Lambda(x,y)}{\partial g_{j}}.$ (53)
@@ B.2 Quantum estimation precision of source distance
$\kappa_{r}=$ $\langle\hat{P}_{r}^{2}\rangle=\eta\sin(2\alpha)+\cos^{2}(\alpha)\kappa_{x}+\sin^{2}(\alpha)\kappa_{y},$
@@
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{{\hbox{N_{{\rm tot}}^{}}}}\left(1-\epsilon^{2}\right)\left(\kappa_{x}\kappa_{y}-\eta^{2}\right)}{\kappa_{r^{\perp}}}.$ (79)

### v2

# Optimal quantum superresolution for full distance between incoherent optical sources in two dimensions

@@ Abstract
The Rayleigh criterion has long served as a fundamental limit for the resolution of optical imaging. Recent advances in multiparameter quantum metrology have led to quantum superresolution that can break this limit and achieve nonvanishing precision in estimating the separation between a pair of closely located incoherent point sources. For two-dimensional optical systems, the quantum superresolution has been studied for the Cartesian components of separation between two incoherent point sources. However, the precision limit of estimating the full distance between two point sources remains unknown so far. In this paper, we investigate the estimation precision of the full distance between two incoherent point sources with arbitrary intensities in a two-dimensional imaging system. Through the multiparameter quantum estimation theory, we obtain the ultimate estimation precision for the distance and show that it remains nonzero when the distance approaches zero, which surpasses the Rayleigh criterion. We further show the dependence of the estimation precision on the relative orientation between the two point sources, which leads to a novel scheme that can enhance the precision by aligning the sources along proper directions if the point-spread functions are not circularly symmetric, and the enhancement is determined by the extent to which the point-spread functions deviate from circular symmetry. Finally, the results are illustrated by incoherent sources with Gaussian point-spread functions.
@@ I Introduction
A notable example is the longstanding challenge of resolving closely spaced optical sources, which is constrained by Rayleigh criterion in classical optics [11]. This criterion sets a diffraction-limited resolution threshold at $d=1.22\lambda/(2{\rm NA})$, with $\lambda$ the wavelength of the light and $\rm NA$} the numerical aperture of the imaging system. To overcome this limit, various approaches have been proposed. One straightforward method is to reduce the wavelength or increase the numerical aperture [12]. More fundamentally, techniques such as microscopy structured illumination [13, 14], photoactivatable localization [15] and stochastic optical reconstruction microscopy [16] have been proposed to surpass the Rayleigh limit by exploiting nonlinear optical responses, stochastic activation of fluorescent emitters, and computational reconstruction algorithms. While these techniques have achieved remarkable progress in imaging resolution, they typically require specialized fluorophores, sophisticated control of activation cycles, or complex image processing [17].
From a statistical perspective, when two incoherent optical sources have significant overlap on the image plane, the precision of estimating their separation is fundamentally limited by the Cramér-Rao lower bound [18, 19]. It follows that the estimation precision of the separation between two optical sources vanishes as the separation approaches zero by direct imaging. This statistical phenomenon is known as Rayleigh’s curse. However, Tsang et al. [20] attacked this problem from the perspective of quantum multiparameter estimation and surprisingly found that the quantum Cramér-Rao bound (QCRB) [21] does not vanish for arbitrarily small separation of two incoherent point sources if the quantum measurements on the point sources are optimized, which addresses Rayleigh’s curse.
@@
In the original superresolution protocol [20], the superresolution was proposed for two optical sources with one-dimensional point-spread functions . However, the point-spread functions of practical optical sources are essentially multidimensional, where the spatial structures of the point-spread functions become crucial to the imaging resolution. In this broader context, multiparameter quantum estimation theory provides a powerful tool to improve the imaging precision [58, 59, 60, 61]. Building upon this foundation, subsequent researches have extended the quantum superresolution to the estimation of Cartesian components of the separation for two-dimensional point sources [62] and for three-dimensional point sources [63, 64, 65], and to more complex imaging tasks such as surface metrology [66], using balanced homodyne detection [67], and estimating the sizes of uniform line and disk sources [68], etc.
In this work, we investigate the quantum precision limits of two-dimensional superresolution for a pair of incoherent point sources with arbitrary relative intensities. Rather than the individual Cartesian components of the separation, we focus on the quantum estimation of the full distance between two point sources. We show that, by optimizing the measurement basis, the quantum Fisher information for the distance remains finite when the two sources get close to each other, which breaks the Rayleigh limit. Furthermore, when the point-spread functions (PSF) of the two point sources are not circularly symmetric, the estimation precision exhibits a nontrivial dependence on the relative orientation between the two point sources. This orientation dependence enables further precision enhancement by aligning the two point sources along specific directions. We optimize the quantum Fisher information of the distance over the azimuth between the two point sources, and obtain the optimal azimuth and the maximum quantum Fisher information for the estimation of source distance. Taking Gaussian PSFs as an example, we find that the optimal orientation corresponds to the minor axis of the PSF.
@@ II.1 Model of optical sources
This question has been addressed for one-dimensional imaging systems by the pioneering work of Tsang et al., and also been considered for the estimation of the orthogonal components of the separation between two point sources in two-dimensional imaging systems, as reviewed in the introduction section. In this work, we will focus on the estimation of the full distance, as well as the relative orientation, between two point sources in two-dimensional imaging systems.
Figure 1: Two point optical sources located at $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$. The distance between the two sources is $r$, and the relative orientation between the two point sources characterized by the azimuth $\alpha$, i.e., the angle between the displacement of the two sources and a given reference axis which is the $x$ axis in the figure.
@@ III.1 Quantum precision limit for separation of point sources
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{$N_{{\rm tot}}$}\left(1-\epsilon^{2}\right)\left(\kappa_{x}\kappa_{y}-\eta^{2}\right)}{\kappa_{r^{\perp}}},$ (20)
@@
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{$N_{{\rm tot}}$}\left(\kappa_{x}\kappa_{y}-\eta^{2}\right)}{\kappa_{r^{\perp}}}.$ (22)
Any deviation from this balance leads to a degradation in precision. Counterintuitively, this suggests that introducing asymmetry to the intensities of two sources, while increasing the visual contrast of the intensity distribution on the image plane, actually reduces the resolvability of the two sources. Physically, this is because the resolution of two optical sources relies on the distinguishability between them and lowering the intensity of either source will decrease its distinguishability from the other. For example, if the intensity of one source is tuned to zero, the visual contrast on the image plane reaches the maximum, but that source can never be distinguished from the other as it is impossible to determine its location when its intensity is zero.
@@ III.2 Optimization for precision limit of distance
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{$N_{{\rm tot}}$}\kappa_{x}\kappa_{y}}{\sin^{2}\alpha\kappa_{x}+\cos^{2}\alpha\kappa_{y}},$ (27)
@@ III.3 Example: two-dimensional Gaussian point-spread functions
Figure 3: Density plot of the distance precision $\mathcal{H}_{r}$ for two Gaussian point-spread functions with respect to the azimuth $\alpha$ and and the ratio $\epsilon$ of the photon number difference to the total photon number as $r\rightarrow 0$. $\sigma_{1}=1$, $\sigma_{2}=1.4$ and $\beta=0.4$. It can be observed that when the azimuth $\alpha$ is fixed, the estimation precision of the distance $r$ will decrease as the ratio $\epsilon$ increases and the optimal estimation precision can be always reached when the photon numbers of the two sources are the same, i.e., $\epsilon=0$, showing the negative impact of photon number difference on the superresolution precision. On the other hand, when $\epsilon$ is fixed, the estimation precision of the distance varies with the azimuth $\alpha$, and the maximum precision of the distance $r$ can be obtained when $\alpha$ is approximately $\pi/7$, which corresponds to the direction along the minor axis of the Gaussian point-spread function, implying the precision of estimating the distance can be enhanced by optimizing the relative orientation between the two point sources.
@@
which implies that it is always possible to increase the superresolution precision by aligning the two point sources whenever the Gaussian point-spread function is not circularly symmetric. On the other hand, when the Gaussian point-spread function has circular symmetry, the overlap between the two point-spread functions always remains the same regardless of the aligning of the two sources, so in this case, the precision becomes orientation-independent and there is no advantage in the precision by aligning the two sources.
@@
To illustrate the above results numerically, Fig. 3 plots the estimation precision $\mathcal{H}_{r}$ with respect to $\alpha$ and $\epsilon$ in the limit $r\rightarrow 0$. It can be seen that for a fixed azimuth $\alpha$, the superresolution precision $\mathcal{H}_{r}$ is maximized at $\epsilon=0$, corresponding to the case of equal photon intensities. As $\epsilon$ increases, $\mathcal{H}_{r}$ decreases and eventually vanishes with strong photon intensity imbalance. Moreover, for a given $\epsilon$, $\mathcal{H}_{r}$ shows clear dependence on the azimuth $\alpha$ between the two point sources, reaching its maximum at
$\alpha_{{\rm opt}}=\tan^{-1}\frac{\sigma_{2}^{2}-\sigma_{1}^{2}-\Re}{2\beta\sigma_{1}\sigma_{2}},$ (47)
which coincides with the minor axis of the Gaussian point-spread function.
IV Conclusion
@@
Moreover, we found that the superresolution precision for distance estimation has dependence on the relative orientation between two point sources, which led us to find that the estimation precision of distance can be significantly enhanced by aligning the two sources along specific directions and derive the optimal relative azimuth and the highest estimation precision of the distance between the two sources. This opens a new avenue for optimizing superresolution strategies in practical systems without circular symmetry.
In summary, our findings refine the current understanding of quantum superresolution and its ultimate capabilities in higher-dimensional imaging systems, offering both theoretical insight and practical guidance for the application of quantum superresolution in realistic scenarios.
@@ Appendix A Resolution limit of direct imaging
$\mathcal{J}_{ij}=\int_{-\infty}^{\infty}dy\int_{-\infty}^{\infty}dx\frac{1}{p_{{\bf g}}(x,y)}\frac{\partial p_{{\bf g}}(x,y)}{\partial g_{i}}\frac{\partial p_{{\bf g}}(x,y)}{\partial g_{j}}.$ (52)
@@ B.2 Quantum estimation precision of source distance
$\kappa_{r}=$ $\langle\hat{P}_{r}^{2}\rangle=\eta\sin(2\alpha)+\cos^{2}\alpha\kappa_{x}+\sin^{2}\alpha\kappa_{y},$
@@
$\left.\mathcal{H}_{r}\right|_{r\rightarrow 0}=\frac{\text{$N_{{\rm tot}}$}\left(1-\epsilon^{2}\right)\left(\kappa_{x}\kappa_{y}-\eta^{2}\right)}{\kappa_{r^{\perp}}}.$ (78)